Grade 12Physical Chemistry

Determine the molecular formula of an oxide of iron in which the mass per cent of iron and oxygen are 69.9 and 30.1 respectively.

sudhanshu
12 Years agoGrade 12

1 Answer

Gaurav
11 Years ago

Elements

Symbol

% by mass

Atomic Mass

% by mass/atomic mass

simple ratio

simple whole number ratio

Iron

Fe

69.9

55.85

69.9/55.85 = 1.25

1.25/1.25=1(divide by smaller number of privious coloumn)

2(multiply by 2 to change in whole number )

Oxygen

O

30.1

16.00

30.1/16.00 =1.88

1.88/1.25=1.5

3


Empirical formula mass of Fe2O3 = 159.7
Molar mass of Fe2O3 = 159.69
So n = 1
Or empirical formula is same as molecular formula